### v2

# Supershear–subshear–supershear rupture sequence during the 2025 Mandalay Earthquake in Myanmar

Abstract
We investigate the rupture dynamics of the Mw 7.7 Mandalay earthquake in Myanmar on March 28, 2025. The earthquake involved a near-vertical strike-slip rupture exceeding 400 km, with shallow slip up to 6 m. A previous inversion result revealed that the rupture initially propagated at supershear velocities ($\sim$6 km/s) near the hypocenter. A unique video recording of surface rupture, captured 124 km south of the epicenter, indicates a deceleration to subshear speeds ($\sim$3 km/s) before reaching the camera location. This deceleration is supported by observed fault-normal acceleration patterns. Satellite imagery also indicated a local minimum in slip (2–3 m) around 50 km south of the epicenter, suggesting a region of reduced stress drop, which likely caused the temporary deceleration. Beyond this point, the rupture appears to have re-established supershear propagation.
1 Introduction
Qualitatively and quantitatively describing coseismic surface displacement, deformation, and rupture in the immediate vicinity of a fault remains a critical challenge for both seismology and earthquake engineering. While strong motion records within a few hundred meters of the fault have been obtained from events such as the 2000 Western Tottori earthquake [7] and the 2023 Kahramanmaraş earthquake [15], these observational opportunities are limited due to the difficulty in predicting rupture locations in advance. Such data significantly improve the spatial resolution of a strong motion distribution and play a crucial role in understanding near-fault ground motion.
At 06:20 UTC on March 28, 2025, a magnitude 7.7 earthquake occurred along the Sagaing fault, which traverses Myanmar longitudinally. The epicenter was located at $22.001^{\circ}$, $95.925^{\circ}$, with a focal depth of 10 km [21]. The event involved the rupture of a near-vertical strike-slip fault spanning approximately 400 km oriented north-south. Severe damage from strong ground motion was reported in Myanmar and Thailand, with numerous instances of surface rupture especially observed in areas along the fault trace.
On May 11, 2025, a video purportedly capturing coseismic surface rupture due to the mainshock was uploaded online (https://www.facebook.com/htin.aung.33/videos/1041579804084512). The footage was taken approximately 124 km south of the epicenter. Based on the posted latitude and longitude information ($20.8821^{\circ}$, $96.0353^{\circ}$), the camera, hereafter referred to as CCTV, was located within a megawatt solar power plant adjacent to the northern part of Thazi. The shadows and timestamp within the video suggested that the camera was facing south.
The video displays a timestamp (in seconds) in the upper right corner, which appears to be about five minutes behind the local standard time when compared to the USGS earthquake origin time. The video begins to subtly shake around 12:46:30 at this timestamp, with shaking intensifying from 33 seconds. Between 35 seconds and 1–2 seconds later, the footage clearly shows right-lateral strike-slip displacement of the ground beyond the fence on the right side of the screen and the gate in the foreground. This recording is likely the world’s first direct observation of surface fault rupture during an earthquake, and it provides a wealth of information from a location where no seismographs were installed.
An initial analysis of the video (https://x.com/Bimaterial/status/1922157655641952512) revealed a pulse-like slip velocity function that completed within approximately two seconds. This result has since been refined through the analysis of numerous additional points within the video, indicating that the maximum slip velocity reached 3 m/s [12], or 4.5 m/s considering field observations [8].
For the Myanmar earthquake, information on near-field ground motion is likely limited to the video analysis described above and the waveform data from the GEOFON station in Naypyidaw (NPW) [11]. This station is located approximately 246 km south of the epicenter and about 2.6 km west of the fault. Consequently, it is crucial to maximize the use of all available data and explore diverse avenues to understand the source behavior of this earthquake.
The rupture primarily propagated southward from the hypocenter, extending over 300 km. Notably, an inversion result suggests the propagation of a supershear rupture at a speed of $\sim 6$ km/s around the hypocenter [10]. Furthermore, in the time-corrected fault-parallel displacement waveforms [11], a rapid displacement at NPW, likely associated with the passage of the rupture front, began approximately 48 seconds after rupture initiation, with no clear S-wave observed prior to this. The average velocity during this interval is $\frac{246}{48}=5.125$ km/s. Even based solely on this record, it is strongly inferred that a significant portion of the rupture propagated at supershear velocities.
Conversely, at the CCTV location, which is closer to the epicenter, the rupture might propagate at subshear rather than supershear speeds. This is supported by [12], who note that the S-wave appears to arrive 1.8 seconds earlier than the rupture front at this site. While the video’s timestamp is inaccurate, which makes the exact arrival times of the S-wave and rupture at the camera location unknown, if a substantial portion of the rupture propagated subshear from the initiation point to the CCTV location, it would contradict the inversion results [10]. This suggests a potential change in rupture propagation velocity.
In this paper, we present results estimating the history of surface strong ground motion and rupture propagation velocity using information from both the video and strong motion records. Specifically, insights from ground motion in the fault-normal direction will be used to demonstrate that while the fault rupture generally propagated at supershear speeds, it likely experienced a deceleration to subshear speeds just before reaching the CCTV location.
2 Methods & Results
2.1 Restriction of rupture front
Herein, the origin time of the earthquake, 2025-03-28 06:20:54 (UTC), is defined as $t=0$ sec. As previously stated, the rupture arrival time at the NPW station is estimated to be $t_{N}=48$ sec. We define $r$ (km) as the epicentral distance, with the CCTV location at $r_{C}=124$ km and the NPW station at $r_{N}=246$ km. Assuming a P-wave velocity of $V_{P}=6.0$ km/s, and given that the rupture cannot propagate faster than this velocity, the rupture front at any given location where $r\leq r_{N}$ must arrive within the time window defined by $V_{P}r\leq t\leq V_{P}r+t_{N}-V_{P}r_{N}$. The lower bound of this inequality signifies that rupture cannot occur before the arrival of the P-wave radiated from the hypocenter. The upper bound accounts for the allowable delay time for the rupture to reach NPW at $t=t_{N}$.
Next, we focus on the observation of a clear S-wave at the CCTV location. To investigate its duration, we analyzed the video footage, which has a resolution of 1920$\times$720 pixels and a frame rate of 30 fps. From each frame of the video, we extracted a 1-pixel-height horizontal line at 200 pixels from the top edge, spanning 1920 pixels in width. These extracted lines were then stacked vertically to construct a composite image (Fig.1).
Figure 1: A composite image by extracting a 1-pixel-height horizontal slice from the video at 200 pixels from the top edge and stacking these slices vertically. The tics and values within the image correspond to the timestamp in seconds displayed in the upper right corner of the original video; the timestamp appears to be approximately five minutes behind the actual local time.
In this image analysis, it is crucial to note that the time reference relies on the seconds indicated by the timestamp in the upper right corner of the video, which differs from time $t$ as defined in this paper. We denote the seconds in the video as $T$. The precise shift between $T$ and $t$ remains unknown.
Focusing on the movement of the sliding gate depicted in Fig.1, subtle shaking can be observed starting around $T=30$ sec. This is likely attributed to the P-waves originating from the hypocenter. We have to note that the exact onset cannot be determined from the image analysis and the P-wave could have arrived before $T=30$. This is because the initial phase amplitude is typically low and we cannot estimate the sensitivity of the image to ground shaking, which inherently depends on the rigidity of the structure holding the camera. Subsequently, intense shaking commences at $T\sim 33$ sec., which is interpreted as the arrival of S-waves. Following this, observations of the right side of the screen indicate that fault slip initiated around $T=35$ sec., with the slip completing within approximately two seconds.
Given the earthquake origin time, the arrival times of waves and rupture at NPW, and the duration of S-wave arrival and slip at CCTV, the hypothetical rupture propagation history is estimated as depicted in Fig.2a. The solid white line represents the direct P-wave radiated at the hypocenter. The broken white line defines the condition for the rupture to reach the NPW station at $t=48$ sec.; the rupture front must always remain between these two bounds. Despite the rupture propagation velocity being nearly equal to $V_{P}$ near the rupture initiation point [10], the observation of an approximately two-second S-wave (indicated by the horizontal black bar in Fig.2a) before the rupture front arrival at the CCTV location suggests that the rupture must have decelerated to about 3.0 km/s prior to reaching CCTV. Considering these factors, it is highly probable that the rupture experienced a temporary deceleration, as depicted in Fig.2a, rather than the constant velocity propagation. While the exact duration of the P-wave at CCTV is unclear, even if the deceleration occurred closer to CCTV, it would not significantly alter the conclusion, given that the P-wave duration in the video is at least three seconds.
Figure 2: (a) A potential rupture scenario illustrating supershear-subshear-supershear propagation from the epicenter to the NPW station (star, 246 km). The subshear rupture speed is 3.0 km/s. The color contour represents the slip pulse velocity with a duration of two seconds. The horizontal black bar indicates the two-second duration of the S-wave observed at the CCTV site (124 km). (b) Slip distribution along the surface fault trace derived from Sentinel-2 imagery (see Supplementary Material for details). The ordinates are common to both (a) and (b)
2.2 Fault-normal acceleration prior to the rupture
The ground motion observed at the CCTV site immediately preceding the onset of rupture suggests subshear rupture propagation. A key observation for that is the arrival of S-wave prior to the rupture [12]. The other one is the rightward displacement of the sliding gate during the approximately two-second interval between the S-wave arrival and the initiation of fault slip as seen in Fig.1. This indicates that the ground experienced substantial acceleration towards the east (left side of the screen), causing the sliding gate to be displaced apparently westward due to inertia.
Given that this is a right-lateral strike-slip fault, the rupture is expected to propagate from the north (right foreground of the screen) to the south (left background). Generally, as such slip propagates, the ground ahead of the rupture front moves towards the left of the screen for subshear rupture and towards the right of the screen for supershear rupture, as shown by 2-D steady-state pulse solutions [5] and 3-D dynamic rupture simulations [3]. However, these models do not account for abrupt rupture deceleration directly before the observation point. Therefore, to fully understand this event, it would be beneficial to interpret the results of kinematic simulations where rupture propagation velocity is controlled.
[…]
We calculated the fault-normal acceleration at a distance of $r=124$ km from the epicenter with 2-D and 3-D models. The 2-D modeling was achieved by employing the integral kernel derived for 2-D velocity fields within the framework of the Boundary Integral Equation Method [19]. The 3-D modeling with a free surface and a planar strike-slip fault down to 20 km depth was done using OpenSWPC, an open-source code utilizing a finite difference method [14]. The methodology on the 2-D and 3-D calculations are detailed in the Supplementary Material. For the fault slip time history, we assumed two distinct slip pulses: uniform supershear rupture with the propagation speed of 5.125 km/s and supershear-subshear-supershear rupture sequence as illustrated in Fig.2b. Both scenarios involve a slip velocity, which is given by $\dot{D}(t)=\sin^{2}{\dfrac{\pi(t-t_{r})}{2}}H(t-t_{r})H(t_{r}+2-t)$ with a duration of two seconds, where $H(\cdot)$ is the Heaviside function. The rupture initiation time, $t_{r}$, is dependent on the epicentral distance, $r$, and follows the scenarios. While the actual duration of fault slip can exceed two seconds at various locations, with some areas peaking more than 10 seconds after rupture initiation [10], the impact of these longer durations would arrive at the CCTV site with an additional lag. Therefore, given that our current analysis focuses only on the period immediately surrounding the rupture’s arrival, these longer slip durations are disregarded.
Our modeling reveals distinct fault-normal acceleration patterns depending on the assumed rupture propagation. In the case of constant supershear rupture (Fig.3a), only a gradually increasing westward acceleration is observed prior to the onset of slip. Conversely, an abrupt change in rupture propagation velocity (Fig.3b) leads to the generation of a significant eastward acceleration commencing two seconds prior to slip initiation. This finding demonstrates robustness, having been reproduced by both 2-D and 3-D simulations. The discrepancies observed after $t\sim 27$ s are attributed to the finite fault extent in the 3-D case, where strong radiation also appears from the bottom edge of the fault. Unlike the constant velocity scenario, Fig.3b also shows discernible ground motion caused by P-waves, which is consistent with the observations from the video analysis. Therefore, these results strongly suggest that the rupture decelerated sharply just before reaching the CCTV location.
3 Discussion
To investigate the reason for the abrupt deceleration of the rupture, we calculated the surface fault slip distribution (Fig.2b) using satellite imagery. North-south surface displacements obtained at 80-m resolution reveal that slip occurred precisely along the known fault trace (Figure S1). By taking their difference between the east and west, we can derive the surface slip distribution. The results show that the slip reached 5 m near the epicenter. However, it exhibits a local minimum of 2–3 m at $\sim 50$ km south of the epicenter. Further south, a widespread slip of 4 m is distributed over an epicentral distance of about 100–250 km (see Supplementary Material for details).
Considering that fault slip distribution is generally proportional to the static stress drop, it can be inferred that a large static stress drop occurred near the epicenter. Conversely, there likely existed a region with a small stress drop, centered $\sim 50$ km from the epicenter ($\sim 21.5^{\circ}$). Therefore, while the rupture accelerated to the P-wave speed near the epicenter due to the large stress drop, the energy release gradually weakened as the rupture propagated around the 50-kilometer point, making it unable to maintain supershear. Theoretically, rupture propagation between the Rayleigh wave speed and $\sqrt{2}$ times the S-wave speed is unstable because the singularity is weaker than the square root; consequently, the energy release rate decreases as the rupture decelerates within this speed range [6]. Hence, the rupture might immediately revert to subshear regime due to the positive feedback between the deceleration of the propagation speed and reduction of energy release rate. Subsequently, beyond the 100-kilometer point, an increase in both stress drop and energy release may have led to the re-establishment of supershear rupture.
The local minimum of the released potency density was also independently identified by a teleseismic inversion analysis [10]. Some structures that may behave as a barrier and reduce the slip at $\sim 21.5^{\circ}$ on the Sagaing fault have been previously suggested. Some investigations have revealed that the location $\sim 21.5^{\circ}$ is a fault segment boundary characterized by the edge of a seismic gap and a discontinuity in the fault dip angle [9, 20]. Our model may thus provide a case study of the interaction between a dramatic change in rupture speed and a geological barrier.
4 Conclusions
This study provides unprecedented insights into the rupture characteristics of the 2025 Mandalay Earthquake (Mw 7.7) along the Sagaing Fault. Through the integrated analysis of a unique coseismic video recording, seismic data, and satellite imagery, we have revealed a dynamic rupture history.
Our findings indicate that the rupture initiated with supershear velocities near the hypocenter but experienced a notable deceleration to subshear speeds as it approached the CCTV observation site. This temporary slowdown is strongly supported by observed fault-normal acceleration patterns and aligns with the presence of a localized minimum in fault slip around 50 km south of the epicenter, suggesting a zone of reduced stress drop. Subsequently, beyond this point, the rupture appears to have re-established its supershear propagation.
This research demonstrates the immense value of direct observations of surface fault rupture, especially in environments where traditional seismic instrumentation is sparse. The detailed rupture dynamics unveiled here contribute significantly to our understanding of complex earthquake behaviors, particularly concerning the factors influencing rupture velocity changes and their implications for near-fault ground motion. Future work will aim to further constrain the precise timing of these velocity transitions and their correlation with geological and stress conditions along the fault.
acknowledgements
This study was supported by ERI JURP 2025-S-B102. This study was conducted using the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, the University of Tokyo. The authors thank Dr. Ryo Okuwaki for their helpful comments.
Data and code availability
The source code for 2-D BEM calculation is available at https://doi.org/10.17605/OSF.IO/MTR76. OpenSWPC, the finite difference code, for our 3-D calculation is available at https://doi.org/10.5281/zenodo.15665702. The surface displacement in Supplementary Material is based on Copernicus Sentinel data (2025) processed by Sentinel Hub (https://dataspace.copernicus.eu/explore-data/data-collections/sentinel-data/sentinel-2).
Competing interests
The authors have no competing interests.
Appendix A Supplementary Material: Methods
A.1 2-D kinematic rupture simulation by a boundary element method
We calculated on-fault acceleration during the rupture propagation by 2-D boundary element method. The numerical integration is nothing but calculating equation (32) of Tada & Madariaga[19]. In their notation, the flat 2-D crack is along the $x_{1}$-axis. We consider strike-slip faulting, where the slip direction and fault-normal direction are along the $x_{1}$- and $x_{2}$-directions, respectively. Hence, our equation is
$\dot{u}_{2}^{i,j,n}=\sum_{k,m}D_{1}^{k,m}\,\dot{K}_{1,u2}^{i,j,k,n,m},$ (S.1)
and we obtain the acceleration waveform by differentiating eq.(S.1) with respect to time.
The slip-rate function is
$D_{1}^{k,m}=\sin^{2}\left(\frac{\pi}{2}\left(t_{m}-t_{r}^{k}\right)\right)\,H\left(t_{m}-t_{r}^{k}\right)\,H\left(t_{r}^{k}+2-t_{m}\right),$ (S.2)
where $H(\cdot)$ is the Heaviside function, and $t_{m}$ is the $m$-th time collocation point. $t_{r}^{k}$ is the rupture arrival time at $x=x_{k}$, where $x_{k}$ is the $k$-th discretized fault segment edge. For Figure 3a and 3b in the main text, we employed the following two scenarios, respectively:
  • If the rupture velocity is 5.125 km/s (constant) everywhere, then $t_{r}^{k}=\dfrac{x_{k}}{5.125}$.
  • If rupture velocity is 3 km/s in 82 km $<x<$ 124 km and is 6 km/s otherwise, then $t_{r}^{k}=\begin{cases}\frac{1}{6}x_{k}&(0\leq x_{k}\leq 82)\\ \frac{82}{6}+\frac{1}{3}(x_{k}-82)&(82<x_{k}<124)\\ \frac{82}{6}+\frac{1}{3}(124-82)+\frac{1}{6}(x_{k}-124)&(124\leq x_{k})\\ \end{cases}$.
In the second scenario, the points where the rupture velocity jumps and the subshear rupture velocity (3 km/s) are chosen such that $t_{r}^{k}=48$ s at $x_{k}=246$ km, satisfying the rupture arrival condition at NPW station.
A.2 3-D kinematic rupture simulation by a finite difference method
The equations of motion and the constitutive relations based on the generalized Zener body were numerically solved using the three-dimensional parallel finite difference code OpenSWPC[14, 13]. A three-dimensional viscoelastic medium with dimensions of 500 km (north-south) $\times$ 200 km (east-west) $\times$ 100 km (vertical) was discretized using a uniform grid with 100-meter spacing in each direction. The medium was assumed to be a homogeneous, half-infinite viscoelastic body with a mass density of 2.7 g/cm3, P-wave velocity of 6.0 km/s, S-wave velocity of 3.5 km/s, and intrinsic attenuation factors of Qp = 600 and Qs = 300. The ground surface was defined at 5 km from the top of the domain, and traction-free boundary conditions[16] were applied there. The earthquake source was represented as a collection of discrete point sources. A rectangular region in the central part of the model, extending 300 km in the north-south direction and 20 km in depth from the ground surface, was designated as the fault plane. Stress changes associated with fault slip[4, 17] were applied to the numerical grid points corresponding to the fault for the duration of the rise time starting from the rupture onset time. The rupture onset time at each point was computed according to the setup described in the main text, with rupture initiation assumed at the northern end of the fault. To suppress artificial reflections of seismic waves from the model boundaries, the ADE-CFS perfectly matched layer (PML) scheme[22] with a thickness of 20 grid points was applied along the outer edges of the computational domain. The model was partitioned into 32 $\times$ 32 segments in the horizontal directions, and the simulation was executed on 256 CPUs. Four MPI processes were launched per CPU, with each process utilizing one-fourth of the CPU’s cores via OpenMP threading. The computation of 60,000 time steps with a time increment of 0.005 s took approximately 30 minutes.
A.3 Estimation of Surface fault slip
To estimate the coseismic surface fault slip, we performed pixel-offset (optical correlation) analysis using Sentinel-2 satellite imagery obtained before and after the earthquake. Here, we used 10 m-resolved Band-3 images of the L2A product, with a central wavelength of 560 nm, obtained on March 20th and 30th, 2025.
For the pixel-offset analysis, we used geoCosiCorr3D [1, 2]. Here, we employed 32 pixels (320 m) as a window size for correlating the images, and estimated northward and eastward displacement for every eight steps (80 m). Estimated surface displacements were mapped in Figure S.4.
The distribution of surface fault slip (Figure 2b in the main text) was estimated from the discrepancies in northward components between both sides of the surface fault trace, assuming that this component corresponds with the fault displacements, as the fault strike is almost north-south, and the east-west component has hardly been detected. Considering the window size of the pixel-offset analysis, we excluded 1 pixel on the fault trace and 2 pixels (160 m) on each side from the fault in the analysis results. We then calculated the average displacements ranging from 3 to 12 pixels (800 m) on each side of the fault on the east-west cross-sections.
Figure S.4: Surface displacement estimated by the pixel-offset (optical correlation) analysis of Sentinel-2 imagery. (a) Northward and (b) Eastward displacements. The red star and the X mark show the locations of the epicenter and CCTV, respectively. Fault traces are based on [18].